OpenStudy (anonymous):

A train travels 360 km at a uniform speed. if speed had been 5 km/hr more, it have taken 1 hour less for the same journey. form the quadratic equation to find the speed of the train.

OpenStudy (anonymous):

if v then d=360 with time t .................................. (1) if (v+5) and d=360 too then time needed is t-1 ............................(2) (1) says: 360=v*t ...............................(3) (2) says: 360=(v+5)*(t-1) ....................(4) Find the quadratic eq from (3) and (4) :)

OpenStudy (anonymous):

so it would be 40 km/h right?

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

yeh because at the end you get two answers- -45 and 40 ... since you cannot have a negative speed it must be 40 XD

OpenStudy (anonymous):

let the speed of the train be x km/hr then time rqd to travel 360 =360/x for the speed of (x+5) km/hr the then time is 360/(x+5) by the problem 360/x - 360/(x+5) =1 hence 1/x - 1/(x+5) =1/360 or 5/(x(x+5)) =1/360 or the required quadratic is x(x+5)=360*5 or x^2 +5x -1800=0 (reqd quadratic) hence x^2 +45x -40x-1800=0 or x(x+45) -40(x+45)=0 or(x+45)(x-40)=0 hence x =40 (neglecting -ve speed of -45 ) thus the required speed of the train is 40 km/hr

OpenStudy (anonymous):

thank u matricked

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

plz help a two digit num is such that the product of the digit is 12. when 36 is added to the number the digits interchange thier places. find quad eq?